OpenStudy (anonymous):

Complete each conversion. Round to the nearest hundredth. 40 mi/h ≈ mi/min

sam (.sam.):

40 miles per hour = 0.666666667 miles per minute

OpenStudy (anonymous):

40mi/h * 1 = 40mi/h 60 mins in an hour, so 1 = h/60min 40mi/h * h/60min = 40mi/h (40/60) mi/min = 40mi/h 40mi/h approx = 0.67 mi/min